Science Report Oct 27th

Crew 214 Science Report 27-Oct-2019

Submitted by Science Officer Andrew Wheeler

1. EVA Suit Maintenance: The hab has been set up following arrival. All backpacks have been plugged in and fully charged. Helmets are clean and contained in their cloth bags except for one helmet, for which, the cloth bag is missing. Radios have been tested and are fully charged. Communication earpieces remain to be tested for faulty performance.

2. Green Hab: The green hab has been activated and is awaiting water and additional soil base before sowing of this seasons crops.

3. Astronomy Dome Operations: No astronomical observations have been planned for this crew rotation.

4. Science Dome Operations: The Science Dome has been activated and is awaiting EVAs for sample acquisition and sample testing.

5. RAM Operations: The RAM Dome has been activated and is ready for contingencies.

6. EVA: No preparations or activities have been undertaken to date. The access routes have been identified to allow sample locations to be identified.

7. Closed Loop Food Waste Study: This study has been initiated and is ongoing. Day’s waste: 79g; consisting of 323 calories, 10g fat, 5g saturated fat, 1179mg sodium, 52g carbohydrates, 7g fibre, 6g sugar, 7g protein.

8. Additional activities: Nil

Copyright © The Mars Society. All rights reserved. | Main Site